the sum of two numbers is 57. If the larger number is increased by 3 it is equal 4 times the smaller number
Found 2 solutions by mananth, ewatrrr:
Answer by mananth(12270)

the sum of two numbers is 57. If the larger number is increased by 3 it is equal 4 times the smaller number
...
if one number is x
the other will be 57-x
..
x+3=4(57-x)
x+3=188-4x
5x=188-3
5x=185
/5
x=37
other number = 57-37=20
...
m.ananth@hotmail.ca

Answer by ewatrrr(10682)


Hi,
Let x represent the larger number. Then the other number is (57-x)
write as we read***
x+ 3 = 4(57-x)
solve
x+ 3 = 228 - 4x
5x = 225
x = 45 the larger number, smaller number is 12
checking our answer
48 = 4*12